The probability that Robin Hood hits a target is 5/6. The probability that Little John hits a target is 1/7. If Robin Hood and Little John each shoot one arrow at the target, what is the probability that they both miss?:
Answer by stanbon(75887)

The probability that Robin Hood hits a target is 5/6. The probability that Little John hits a target is 1/7. If Robin Hood and Little John each shoot one arrow at the target, what is the probability that they both miss?:
---------------
P(Robin Hood misses) = 1/6
P(Little John misses) = 6/7
------------------------------
The events are independent, so
P(RH misses and LJ misses) = (1/6)(6/7) = 1/7
------------------
Cheers,
Stan H.